Skip to content

OSG Technology Area Meeting, 4 June 2018

Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Carl, Derek, Edgar, Mat, Suchandra, TimT

Announcements

GitHub acquired by Microsoft! This shouldn't have any effect in the short term.

Triage Duty

  • This week: Suchandra
  • Next week: BrianL
  • 16 (-1) open tickets
  • 3 (-1) open LCMAPS VOMS tickets

JIRA

# of tickets Δ State
113 +4 Open
30 -1 In Progress
6 -1 Ready for Testing
4 -4 Ready for Release

OSG Software Team

  • Support:
    • We will discuss Freshdesk workflows during the next doc focus afternoon
    • Ticket publicity is going to be discussed: for the time being if someone other than the ticket owner asks for ticket history, feel free to send it to them
  • Mail not being sent to the OSG software mailing list (mail eventually received, checking with FNAL to see if there were known issues on their end)
  • JIRA + GitHub integration may no longer work (free trial for the paid plugin expired)
  • OASIS login node needs its WN tarball install updated regularly (puppet?)
  • Office hours: Tuesdays 2-3pm CDT

Discussion

  • AI (Mat) - Create a freshdesk ticket for OASIS login
  • AI (TimT) - Create a freshdesk ticket for a ticket that wasn't created from mail sent to [email protected]

Support Update

  • User certificates (TimT): handled many renewal requests. Most common issue was that people were not logged in, and we could have addressed this in subsequent emails/documentation. Additionally, next time we should set the "Reply-To" to [email protected] so that TimT doesn't have to field all support requests.
  • UConn (Suchandra) - Richard has allocations on various clusters and wants to use them for GlueX jobs, wants to reduce number of CEs necessary
  • UConn (Edgar): Lowered pilot lifetime since jobs opportunistic jobs were taking over the cluster dedicated to GlueX
  • CHTC (Edgar) - no GLOW pilots running due to pilot certs expiring. Fixed now; certs were requested but not put in the proper location.
  • Central collector (Derek) - FIU TCP issues blocking connections to the central collector. Proven by speed of light. Continuing to poke sites this week

OSG Release Team

3.4.13 Δ Status
12 +1 Open
9 +0 In Progress
5 +1 Ready for Testing
4 +4 Ready for Release
30 +6 Total
  • OSG 3.4.13
    • Testing
      • HTCondor 8.6.11
      • HTCondor 8.7.8
      • Drop consider_as_osg form *.repo files
      • OWAMP 3.5.6
    • Ready for Release
      • voms-proxy-direct
      • Pegasus 4.8.2
      • Singularity 2.5.1
      • CMVFS 2.5.0
  • Data
    • Nothing yet
  • GOC
    • Testing
      • repo-update-cadist: RPM packaging
    • Ready for Release
      • Nothing yet

Discussions

  • AI (TimT): update SW release doc for post-service migrations (SOFTWARE-3276)
  • AI (TimT): copy data release doc to create a basic GOC release doc

OSG Investigations Team

GOC Transitions

Done with transitions!?!

We are all but done with the transition of services from IU. There are still a few items on our list:

  • (Ongoing) Getting sites to transition to the new HTCondor collector
  • Minor hiccup with OSG-WN tarball in OASIS home directory. Still investigating.

And a few longer term goals:

  • Most services are in puppet. Just need to put the finishing touches on it. For example, the changes made in the last week with all of the OIM, Ticket,... redirections.
  • Most services are in check_mk (active monitoring), but need to add specific tests. For example, we check that HTTP(s) is responding to requests, but we would like to make sure it is responding with correct info.

Presentations

Gave a StashCache presentation at the GPN annual meeting. Went over very well. Highlighted the science drivers for StashCache.

Ongoing

Discussions

AI (BrianL, Derek) - Investigate mailing list outage (possibly related to DNS security changes made last week) and announce outage if necessary